OpenStudy (anonymous):

Solve for x in the following equation: -8.9 + x = 28.48 A. x = -3.2 B. x = 3.2 C. x = 37.38 D. x = 19.58

pooja195 (pooja195):

Start by adding 8.9 to both sides 28.48+8.9=?

OpenStudy (anonymous):

start, and end

OpenStudy (anonymous):

ok

pooja195 (pooja195):

Aight. -8.9+8.9=? 28.48+8.9=? :P

OpenStudy (anonymous):

-8.9+8.9 = 0 :P

pooja195 (pooja195):

Good and what about the other one? :P

OpenStudy (anonymous):

28.48 + 08.90 = 37 .38

OpenStudy (skullpatrol):

You are done :-)

OpenStudy (anonymous):

thank chu!

OpenStudy (skullpatrol):

Thanks for asking :D